OpenStudy (anonymous):

Simplify the following quotient of complex numbers into the form a + bi. 1+8i ------- -6+i

OpenStudy (anonymous):

multiply numerator and denominator by the conjugate of the denominator. the conjugate of \[a+bi\] is \[a-bi\] and \[(a+bi)(a-bi)=a^2+b^2\] so you will get a real number in the denominator

OpenStudy (anonymous):

in this case you get \[\frac{(1+8i)(-6-i)}{6^2+1^2}=\frac{(1+8i)(-6-i)}{37}\]

OpenStudy (anonymous):

so your actual job it to multiply out in the numerator. is it clear how to do that?

OpenStudy (anonymous):

in think so

OpenStudy (anonymous):

yeah i get it

OpenStudy (anonymous):

then don't forget to split the number up so it is in standard form

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

and let me know when you get \[\frac{2}{37}-\frac{49}{37}i\]

OpenStudy (anonymous):

i think its 47/37